OpenStudy (gabylovesyou):

In the below system, solve for y in the first equation. x − 3y = −6 2x − 7y = 10 A. one thirdx + 2 B. negative one thirdx + 6 C. −x + 2 D. negative one thirdx + 2

OpenStudy (bemoseemo):

just need a walkthrough or can I just this away?

OpenStudy (gabylovesyou):

Wym?

OpenStudy (gabylovesyou):

I have no idea how to do this .. Lmfao 😂

OpenStudy (nitroster):

I know how to

OpenStudy (bemoseemo):

can I say the answer?

OpenStudy (nitroster):

stick with bemo... :P But bemo, explain it. She doesnt know it..

OpenStudy (gabylovesyou):

Can you try to explain? :)

OpenStudy (bemoseemo):

ill try to explain as well as I can

OpenStudy (gabylovesyou):

Okay thank you.

OpenStudy (bemoseemo):

so first we have x -3y, to help find y, we also have 2x -7y

OpenStudy (gabylovesyou):

Ok

OpenStudy (bemoseemo):

try to finish this: x−3y=−6 (x−3y)+(−x)=−6+(−x) x−3y−x=−6−x

OpenStudy (gabylovesyou):

3y = -6-x ?

OpenStudy (bemoseemo):

this is a screenshot of all the work explained out

OpenStudy (hamde):

x=3y-6 2(3y-6)-7y=10 6y-12-7y=10 -y-12=10 -y=22 y=-22 then x=3*-22-6 x=-66-6 x=-72

OpenStudy (bemoseemo):

one thirdx + 2 once you simplify it

OpenStudy (gabylovesyou):

Oh ok

OpenStudy (gabylovesyou):

And now for the other one?

OpenStudy (bemoseemo):

I have to go, but try to figure out how I worked it through in that screenshot k?

OpenStudy (gabylovesyou):

Ok thank you! :3